what is the molarity of an acetic acid solution that requires 40.0 ml of 0.100 M NaOH to titrate a 20.0 ml sample to the end point?
1
Expert's answer
2018-05-30T05:55:39-0400
CH3COOH + NaOH = CH3COONa + H2O
Substances react in equal equivalent quantities: n (NaOH) = n (CH3COOH)

C = n/V
n = C*V
C(NaOH)*V(NaOH) = C(CH3COOH)*V(CH3COOH)
C(CH3COOH) = C(NaOH)*V(NaOH) / V(CH3COOH) = 0.1*40 / 20 = 0.2

Answer: CM (CH3COOH) = 0.2 M